# New versions of Frobenius and integral closure of ideals

### v2

@@ 4. Adjusted Ideal Closures in Semigroup Rings
We now consider the class of rings $k[S]$, where $k$ is a field and $S$ is a submonoid of the free abelian group $M=\mathbb{Z}^{d}$ of rank $d$. We have that $M$ naturally sits inside $V=M\otimes_{\mathbb{Z}}\mathbb{R}\cong\mathbb{R}^{d}$ and we can utilize convex geometry in $\mathbb{R}^{d}$ to analyze monomial ideals in $k[S]$. We will briefly describe the tools we need; we refer the reader to books such as [Ful93] or [CLS11] for further background on affine semigroup rings.
Setting 4.1.
Fix a field $k$. Let $M$ be a rank $d$ free abelian group, $S$ a finitely generated submonoid of $M$ such that $S$ generates $M$ as a group, and $V=M\otimes_{\mathbb{Z}}\mathbb{R}$. Let $\sigma^{\vee}=\mathrm{cone}(S)\subset V$, and $\sigma=\mathrm{cone}(\mathbf{c}_{1},\ldots,\mathbf{c}_{n})\subset V^{*}$ the dual cone to $\sigma^{\vee}$. That is, $\mathbf{v}\in\sigma^{\vee}$ if and only if $\mathbf{v}\in V$ and $\langle\mathbf{c}_{i},\mathbf{v}\rangle\geq 0$ for all $i=1,\ldots,n$. Let $\bar{S}=\sigma^{\vee}\cap M$. Thus, $\bar{S}$ is the saturation of $S$ inside $M$, and $\bar{R}=k[\bar{S}]$ is the integral closure of $R=k[S]$ inside $k[M]$. Note that for any ideal $J$ of $R$, $(\operatorname{Exp}(J)+\sigma^{\vee})\cap M=\operatorname{Exp}(J\bar{R})$.
Let $I$ and $\mathfrak{a}$ be monomial ideals of $R$, so that there exist finite sets $\mathcal{A},\mathcal{B}\subset S$ such that $\mathfrak{a}=(x^{\mathbf{a}}\mid\mathbf{a}\in\mathcal{A})$ and $I=(x^{\mathbf{b}}\mid\mathbf{b}\in\mathcal{B})$. Let $A=(\mathbf{a}_{1}|\cdots|\mathbf{a}_{r})$, and let $\Delta=\{\mathbf{u}\in\mathbb{R}_{\geq 0}^{r}\mid|\mathbf{u}|_{1}=1\}$. Let $H=\mathrm{Hull}(\mathbf{a}_{1},\ldots,\mathbf{a}_{r})$ be the convex hull of the columns of $A$, and notice that $H=\{A\mathbf{u}\mid\mathbf{u}\in\Delta\}$.
@@
Given 4.1, for each $n\in\mathbb{Z}_{\geq 0}$ and $s\in\mathbb{R}_{>0}$, we have
@@
    $sH\subset\operatorname{Exp}(\mathfrak{a}^{\lfloor s-d\rfloor})+\sigma^{\vee}$.
@@
    For any $n\in\mathbb{N}$, $\mathfrak{a}^{n}$ is generated by monomials of the form $A\mathbf{w}$, where $\mathbf{w}\in\mathbb{N}^{r}$ and $|\mathbf{w}|_{1}=n$. Letting $\mathbf{u}=\frac{1}{n}\mathbf{w}$, we have that $\mathbf{u}\in\Delta$ and so $A\mathbf{w}=nA\mathbf{u}\in nH$. Therefore $\operatorname{Exp}(\mathfrak{a}^{n})\subset nH+\sigma^{\vee}.$
@@
    On the other hand, if $\mathbf{v}\in sH$ and $s>0$, then there exists $\mathbf{u}\in\Delta$ such that $\mathbf{v}=sA\mathbf{u}$. Furthermore, by Carathéodory’s theorem, we may assume that at most $d+1$ of the components of $u$ are nonzero. Let $\mathbf{w}\in\mathbb{N}^{r}$ be given by $w_{i}=\lfloor su_{i}\rfloor$. Now,
@@
    Since $|\mathbf{w}|_{1}=\sum_{i}\lfloor su_{i}\rfloor>\sum_{i:u_{i}\neq 0}(su_{i}-1)\geq s-d-1$, we have that $|\mathbf{w}|_{1}\geq\lfloor s-d\rfloor$ and so $A\mathbf{w}\in\operatorname{Exp}(\mathfrak{a}^{\lfloor s-d\rfloor})$. Thus, we have that $sH\subset\operatorname{Exp}(\mathfrak{a}^{\lfloor s-d\rfloor})+\sigma^{\vee}$, and so $sH\cap\bar{S}\subset\operatorname{Exp}(\mathfrak{a}^{\lfloor s-d\rfloor}\bar{R})$.
@@ 4.1. Joint Hilbert-Kunz multiplicity
In [Vra08], the author defined the joint Hilbert-Kunz multiplicity, $e_{HK}(\mathfrak{a}^{t};I),$ of two finite colength ideals $I,\mathfrak{a}$ in a Noetherian local ring $(R,\mathfrak{m})$ of dimension $d\geq 1.$ Letting $t>0$ be a fixed real number,
$e_{HK}(\mathfrak{a}^{t};I)=\lim\limits_{q\rightarrow\infty}\frac{1}{q^{d}}\ell\left(\frac{R}{\mathfrak{a}^{\lceil tq\rceil}I^{[q]}}\right).$
It is known that the Hilbert-Kunz multiplicity of finite length ideals in a semigroup ring can be computed as the volume of a finite union of rational polytopes. The following result proves that the same holds for the joint Hilbert-Kunz multiplicity.
@@
Given 4.1, suppose $k$ is of prime characteristic $p>0$, and let $I$ and $\mathfrak{a}$ be ideals in $R$ of finite colength, and let $t>0$ be a fixed real number. If we set
@@
First note that it is enough to prove the statement when $R$ is a normal ring, as if $R$ is not normal, then from [Vra08, Lemma 3.3] we get that $e_{HK}(\mathfrak{a}^{t};I)=e_{HK}(\mathfrak{a}^{t}\bar{R};I\bar{R})$ as $\overline{R}$ is a module-finite extension of $R$.
@@
is a finite union of rational polytopes, say $P_{t,1},\ldots,P_{t,\beta},$ containing the origin such that for all $i\neq j,$ $P_{t,i}\cap P_{t,j}$ is a rational polytope of dimension $<d.$ Using the same arguments as in the proof of [Eto02, Theorem 2.2], we get
(4.1) $\operatorname{vol}\overline{\mathcal{P}_{t}}=\sum_{i=1}^{\beta}\operatorname{vol}P_{t,i}=\sum_{i=1}^{\beta}\lim\limits_{q\rightarrow\infty}\frac{\mid qP_{t,i}\cap\mathbb{Z}^{d}\mid}{q^{d}}=\lim\limits_{q\rightarrow\infty}\frac{\mid q\mathcal{P}_{t}\cap\mathbb{Z}^{d}\mid}{q^{d}}.$
In order to compute the joint Hilbert-Kunz multiplicity, we observe that if $x^{\mathbf{v}}\in\mathfrak{a}^{\lceil tq\rceil}I^{[q]}$, then $\mathbf{v}\in A\mathbf{w}+q\mathcal{B}+\sigma^{\vee}$, where $\mathbf{w}\in\mathbb{N}^{r}$ and $|\mathbf{w}|_{1}=\lceil tq\rceil$, so we have
$\mathbf{v}\in\lceil tq\rceil H+q\mathcal{B}+\sigma^{\vee}\subset q(tH+\mathcal{B}+\sigma^{\vee})=\sigma^{\vee}\setminus q\mathcal{P}_{t}.$
Therefore, if $\mathbf{v}\in q\mathcal{P}_{t}$, then $x^{\mathbf{v}}\notin\mathfrak{a}^{\lceil tq\rceil}I^{[q]}$, and so $\ell(R/\mathfrak{a}^{\lceil tq\rceil}I^{[q]})\geq|q\mathcal{P}_{t}\cap\mathbb{Z}^{d}|$.
On the other hand, suppose that $s>t$, and $q$ is large enough that $\lfloor sq-d\rfloor\geq\lceil tq\rceil$. If $\mathbf{v}\in S\setminus q\mathcal{P}_{s}$, then
$\mathbf{v}\in q\mathcal{B}+sqH+\sigma^{\vee}\subset$ $\operatorname{Exp}(I^{[q]})+\operatorname{Exp}(\mathfrak{a}^{\lfloor sq-d\rfloor})+\sigma^{\vee}$
$=$ $\operatorname{Exp}(\mathfrak{a}^{\lfloor sq-d\rfloor}I^{[q]})+\sigma^{\vee}\subset\operatorname{Exp}(\mathfrak{a}^{\lceil tq\rceil}I^{[q]})+\sigma^{\vee}.$
Therefore $x^{\mathbf{v}}\in\mathfrak{a}^{\lceil tq\rceil}I^{[q]}$. Hence $\ell(R/\mathfrak{a}^{\lceil tq\rceil}I^{[q]})\leq|q\mathcal{P}_{s}\cap\mathbb{Z}^{d}|$.
Thus, we have that, for any $s>t$ and $q\gg 0$,
$|q\mathcal{P}_{t}\cap\mathbb{Z}^{d}|\leq\ell(R/\mathfrak{a}^{\lceil tq\rceil}I^{[q]})\leq|q\mathcal{P}_{s}\cap\mathbb{Z}^{d}|$
@@
$\lim_{q\to\infty}\frac{|q\mathcal{P}_{t}\cap\mathbb{Z}^{d}|}{q^{d}}\leq e_{HK}(\mathfrak{a}^{t};I)\leq\lim_{q\to\infty}\frac{|q\mathcal{P}_{s}\cap\mathbb{Z}^{d}|}{q^{d}}.$
@@ 4.2. The adjusted closures in affine semigroup rings
In this subsection we show that the calculation of the adjusted ideal closures in affine semigroup rings are computable in terms of the convex geometry of the associated exponent sets.
@@
Given 4.1, suppose $k$ is of prime characteristic $p>0$, and let $t>0$ and $x^{\mathbf{v}}\in R$ be a monomial.
@@
Suppose that there exists $x^{\mathbf{c}}\in R$, $s\geq 0$ such that for $q\gg 0$, $x^{\mathbf{c}}x^{q\mathbf{v}}\mathfrak{a}^{\lceil tq\rceil}\subset I^{[q]}\mathfrak{a}^{\lceil sq\rceil}$. That is,
$\mathbf{c}+q\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tq\rceil})\subset q\mathcal{B}+\operatorname{Exp}(\mathfrak{a}^{\lceil sq\rceil})+\sigma^{\vee}.$
@@
$\mathbf{c}+q\mathbf{v}+\lceil tq+d\rceil H$ $\subset\mathbf{c}+q\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tq\rceil})+\sigma^{\vee}$
$\subset q\mathcal{B}+\operatorname{Exp}(\mathfrak{a}^{\lceil sq\rceil})+\sigma^{\vee}$
$\subset q\mathcal{B}+\lceil sq\rceil H+\sigma^{\vee}.$
Since $H\subset\sigma^{\vee}$, $\lceil sq\rceil H+\sigma^{\vee}\subset sqH+\sigma^{\vee}$. Therefore, dividing through by $q$, we have that for $q\gg 0$,
$\frac{1}{q}\mathbf{c}+\mathbf{v}+\frac{\lceil tq+d\rceil}{q}H\subset\mathcal{B}+sH+\sigma^{\vee}.$
Now, since $\mathcal{B}+sH+\sigma^{\vee}$ is a closed set, for any $\mathbf{h}\in H$ and $q\geq 1$,
$\mathbf{v}+t\mathbf{h}=\lim_{q\to\infty}\left(\frac{1}{q}\mathbf{c}+\mathbf{v}+\frac{\lceil tq+d\rceil}{q}\mathbf{h}\right)\in\mathcal{B}+sH+\sigma^{\vee}.$
@@
Given 4.1, suppose $k$ is of prime characteristic $p>0$, and let $t>0$ and $x^{\mathbf{v}}\in R$ be a monomial.
@@
    If $\mathbf{v}+tH\subset\mathcal{B}+\sigma^{\vee}$, then $x^{\mathbf{v}}\in(I\bar{R})^{F\mathfrak{a}^{t}}$ and $x^{\mathbf{v}}\in(I\bar{R})^{*\mathfrak{a}^{t}}$.
@@
    If $\mathbf{v}+tH\subset\mathcal{B}+tH+\sigma^{\vee}$, then $x^{\mathbf{v}}\in{{}^{\mathfrak{a}^{t}}(I\bar{R})^{*}}$.
@@
    If $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ for some $s>t$, then $x^{\mathbf{v}}\in{}^{\mathfrak{a}^{t}}(I\bar{R})^{F}$.

### v3

@@ 6. Adjusted ideal closures in semigroup rings
We now consider the class of rings $k[S]$, where $k$ is a field and $S$ is a submonoid of the free abelian group $M=\mathbb{Z}^{d}$ of rank $d$. We have that $M$ naturally sits inside $V=M\otimes_{\mathbb{Z}}\mathbb{R}\cong\mathbb{R}^{d}$ and we can utilize convex geometry in $\mathbb{R}^{d}$ to analyze monomial ideals in $k[S]$. We will briefly describe the objects we need; we refer the reader to books such as [Ful93] or [CLS11] for further background on affine semigroup rings.
Setting 6.1.
Fix a field $k$. Let $M$ be a rank $d$ free abelian group, $S$ a finitely generated submonoid of $M$ such that $S$ generates $M$ as a group, and $V=M\otimes_{\mathbb{Z}}\mathbb{R}$. Let $\sigma^{\vee}=\mathrm{cone}(S)\subset V$, and $\sigma=\mathrm{cone}(\mathbf{c}_{1},\ldots,\mathbf{c}_{n})\subset V^{*}$ the dual cone to $\sigma^{\vee}$. That is, $\mathbf{v}\in\sigma^{\vee}$ if and only if $\mathbf{v}\in V$ and $\langle\mathbf{c}_{i},\mathbf{v}\rangle\geq 0$ for all $i=1,\ldots,n$. Let $\overline{S}=\sigma^{\vee}\cap M$. Thus, $\overline{S}$ is the saturation of $S$ inside $M$, and $\overline{R}=k[\overline{S}]$ is the integral closure of $R=k[S]$ inside $k[M]$. Given a monomial ideal $I\subset k[M]$, we will write $\operatorname{Exp}I\subset M$ for the set of exponents appearing in monomials in $I$, that is, $\operatorname{Exp}I=\{\mathbf{v}\in S\mid x^{\mathbf{v}}\in I\}$. Note that for any ideal $J$ of $R$, $(\operatorname{Exp}(J)+\sigma^{\vee})\cap M=\operatorname{Exp}(J\overline{R})$.
Let $I$ and $\mathfrak{a}$ be monomial ideals of $R$, so that there exist finite sets $\mathcal{A},\mathcal{B}\subset S$ such that $\mathfrak{a}=(x^{\mathbf{a}}\mid\mathbf{a}\in\mathcal{A})$ and $I=(x^{\mathbf{b}}\mid\mathbf{b}\in\mathcal{B})$. Let $A=(\mathbf{a}_{1}|\cdots|\mathbf{a}_{r})$, and let $\Delta=\{\mathbf{u}\in\mathbb{R}_{\geq 0}^{r}\mid|\mathbf{u}|_{1}=1\}$, where $|\mathbf{u}|_{1}$ is the sum of the components of $\mathbf{u}$. Let $H=\mathrm{Hull}(\mathbf{a}_{1},\ldots,\mathbf{a}_{r})$ be the convex hull of the columns of $A$, and notice that $H=\{A\mathbf{u}\mid\mathbf{u}\in\Delta\}$.
@@
Given 6.1, for each $n\in\mathbb{Z}_{\geq 0}$ and $s\in\mathbb{R}_{>0}$, we have
@@
    $sH\subset\operatorname{Exp}(\mathfrak{a}^{\lfloor s-d\rfloor})+\sigma^{\vee}$ and $sH\cap\overline{S}\subseteq\operatorname{Exp}(\mathfrak{a}^{\lfloor s-d\rfloor}\overline{R})$.
@@
    The ideal $\mathfrak{a}^{n}$ is generated by monomials of the form $x^{A\mathbf{w}}$, where $\mathbf{w}\in\mathbb{N}^{r}$ and $|\mathbf{w}|_{1}=n$. Letting $\mathbf{u}=\frac{1}{n}\mathbf{w}$, we have that $\mathbf{u}\in\Delta$ and so $A\mathbf{w}=nA\mathbf{u}\in nH$. Therefore $\operatorname{Exp}(\mathfrak{a}^{n})\subset nH+\sigma^{\vee}.$
@@
    On the other hand, if $\mathbf{v}\in sH$ and $s>0$, then there exists $\mathbf{u}\in\Delta$ such that $\mathbf{v}=sA\mathbf{u}$. Furthermore, by Carathéodory’s theorem, we may assume that at most $d+1$ of the components of $\mathbf{u}$ are nonzero. Let $\mathbf{w}\in\mathbb{N}^{r}$ be given by $w_{i}=\lfloor su_{i}\rfloor$. Thus each component of $s\mathbf{u}-\mathbf{w}$ is nonnegative, and so $A(s\mathbf{u}-\mathbf{w})$ is a nonnegative linear combinations of the vectors $\mathbf{a}_{i}$. Since each $\mathbf{a}_{i}$ is in $\sigma^{\vee}$, so is $A(s\mathbf{u}-\mathbf{w})$. Hence
@@
    Since $|\mathbf{w}|_{1}=\sum_{i}\lfloor su_{i}\rfloor>\sum_{i:u_{i}\neq 0}(su_{i}-1)\geq s-d-1$, we have that $|\mathbf{w}|_{1}\geq\lfloor s-d\rfloor$ and so $A\mathbf{w}\in\operatorname{Exp}(\mathfrak{a}^{\lfloor s-d\rfloor})$. Thus, we have that $sH\subset\operatorname{Exp}(\mathfrak{a}^{\lfloor s-d\rfloor})+\sigma^{\vee}$, and so $sH\cap\overline{S}\subset\operatorname{Exp}(\mathfrak{a}^{\lfloor s-d\rfloor}\overline{R})$.
@@ 6.1. Joint Hilbert-Kunz multiplicity
In [Vra08], the author defined the joint Hilbert-Kunz multiplicity, $e_{HK}(\mathfrak{a}^{t};I),$ of two finite colength ideals $I,\mathfrak{a}$ in a local ring $(R,\mathfrak{m})$ of dimension $d\geq 1.$ Letting $t\in\mathbb{R}_{>0}$ be a fixed real number,
$e_{HK}(\mathfrak{a}^{t};I)=\lim\limits_{e\rightarrow\infty}\frac{1}{p^{ed}}\lambda_{R}\left(\frac{R}{\mathfrak{a}^{\lceil tp^{e}\rceil}I^{[p^{e}]}}\right).$
It is known that the Hilbert-Kunz multiplicity of finite colength ideals in a semigroup ring can be computed as the volume of a finite union of rational polytopes due to work of Watanabe [Wat00] and Eto [Eto02]. The following result proves that the same holds for the joint Hilbert-Kunz multiplicity.
@@
Given 6.1, suppose $k$ is of prime characteristic $p>0$, and let $I$ and $\mathfrak{a}$ be ideals in $R$ of finite colength, and let $t\in\mathbb{R}_{>0}$. If we set
@@
First note that it is enough to prove the statement when $R$ is a normal ring. Indeed, if $R$ is not normal, then observe that $\overline{R}$ is a finite $R$-module and taking $W=R\setminus\{0\},$ we have $W^{-1}R=W^{-1}\overline{R}=\operatorname{Frac}(R)$ implying that $e_{HK}(\mathfrak{a}^{t};I)=e_{HK}(\mathfrak{a}^{t}\overline{R};I\overline{R})$ (see [Vra08, Lemma 3.3]).
@@
is a finite union of rational polytopes, say $P_{t,1},\ldots,P_{t,\beta},$ containing the origin such that for all $i\neq j,$ $P_{t,i}\cap P_{t,j}$ is a rational polytope of dimension <d.<d. Using the same arguments as in the proof of [Eto02, Theorem 2.2], we get
(6.1) $\operatorname{vol}\overline{\mathcal{P}_{t}}=\sum_{i=1}^{\beta}\operatorname{vol}P_{t,i}=\sum_{i=1}^{\beta}\lim\limits_{p^{e}\rightarrow\infty}\frac{\mid p^{e}P_{t,i}\cap\mathbb{Z}^{d}\mid}{p^{ed}}=\lim\limits_{e\rightarrow\infty}\frac{\mid p^{e}\mathcal{P}_{t}\cap\mathbb{Z}^{d}\mid}{p^{ed}}.$
In order to compute the joint Hilbert-Kunz multiplicity, we observe that if $x^{\mathbf{v}}\in\mathfrak{a}^{\lceil tp^{e}\rceil}I^{[p^{e}]}$, then $\mathbf{v}\in A\mathbf{w}+p^{e}\mathcal{B}+\sigma^{\vee}$, where $\mathbf{w}\in\mathbb{N}^{r}$ and $|\mathbf{w}|_{1}=\lceil tp^{e}\rceil$, so we have
$\mathbf{v}\in\lceil tp^{e}\rceil H+p^{e}\mathcal{B}+\sigma^{\vee}\subset p^{e}(tH+\mathcal{B}+\sigma^{\vee})=\sigma^{\vee}\setminus p^{e}\mathcal{P}_{t}.$
Therefore, if $\mathbf{v}\in p^{e}\mathcal{P}_{t}$, then $x^{\mathbf{v}}\notin\mathfrak{a}^{\lceil tp^{e}\rceil}I^{[p^{e}]}$, and so $\lambda_{R}(R/\mathfrak{a}^{\lceil tp^{e}\rceil}I^{[p^{e}]})\geq|p^{e}\mathcal{P}_{t}\cap\mathbb{Z}^{d}|$.
On the other hand, suppose that $s>t$, and $p^{e}$ is large enough that $\lfloor sp^{e}-d\rfloor\geq\lceil tp^{e}\rceil$. If $\mathbf{v}\in S\setminus p^{e}\mathcal{P}_{s}$, then
$\mathbf{v}\in p^{e}\mathcal{B}+sp^{e}H+\sigma^{\vee}\subset$ $\operatorname{Exp}(I^{[p^{e}]})+\operatorname{Exp}(\mathfrak{a}^{\lfloor sp^{e}-d\rfloor})+\sigma^{\vee}$
$=$ $\operatorname{Exp}(\mathfrak{a}^{\lfloor sp^{e}-d\rfloor}I^{[p^{e}]})+\sigma^{\vee}\subset\operatorname{Exp}(\mathfrak{a}^{\lceil tp^{e}\rceil}I^{[p^{e}]})+\sigma^{\vee}.$
Therefore $x^{\mathbf{v}}\in\mathfrak{a}^{\lceil tp^{e}\rceil}I^{[p^{e}]}$. Hence $\lambda_{R}(R/\mathfrak{a}^{\lceil tp^{e}\rceil}I^{[p^{e}]})\leq|p^{e}\mathcal{P}_{s}\cap\mathbb{Z}^{d}|$.
Thus, we have that, for any $s>t$ and $e\gg 0$,
$|p^{e}\mathcal{P}_{t}\cap\mathbb{Z}^{d}|\leq\lambda_{R}(R/\mathfrak{a}^{\lceil tp^{e}\rceil}I^{[p^{e}]})\leq|p^{e}\mathcal{P}_{s}\cap\mathbb{Z}^{d}|$
@@
$\lim_{e\to\infty}\frac{|p^{e}\mathcal{P}_{t}\cap\mathbb{Z}^{d}|}{p^{ed}}\leq e_{HK}(\mathfrak{a}^{t};I)\leq\lim_{e\to\infty}\frac{|p^{e}\mathcal{P}_{s}\cap\mathbb{Z}^{d}|}{p^{ed}}.$
@@ 6.2. The adjusted closures in affine semigroup rings
In this subsection we show that the calculation of the adjusted closure operations in affine semigroup rings are computable in terms of the convex geometry of the associated exponent sets. This is well-known in the setting of usual integral closure, see for example [SH06, Prop. 1.4.6].
@@
Given 6.1, suppose $k$ is of prime characteristic $p>0$, and let $t\in\mathbb{R}_{>0}$ and $x^{\mathbf{v}}\in R$ be a monomial.
@@
Suppose that there exists $x^{\mathbf{c}}\in R$, $s\geq 0$ such that for $p^{e}\gg 0$, $x^{\mathbf{c}}x^{p^{e}\mathbf{v}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}\mathfrak{a}^{\lceil sp^{e}\rceil}$. That is,
$\mathbf{c}+p^{e}\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tp^{e}\rceil})\subset p^{e}\mathcal{B}+\operatorname{Exp}(\mathfrak{a}^{\lceil sp^{e}\rceil})+\sigma^{\vee}.$
@@
$\mathbf{c}+p^{e}\mathbf{v}+\lceil tp^{e}+d\rceil H$ $\subset\mathbf{c}+p^{e}\mathbf{v}+\operatorname{Exp}(\mathfrak{a}^{\lceil tp^{e}\rceil})+\sigma^{\vee}$
$\subset p^{e}\mathcal{B}+\operatorname{Exp}(\mathfrak{a}^{\lceil sp^{e}\rceil})+\sigma^{\vee}$
$\subset p^{e}\mathcal{B}+\lceil sp^{e}\rceil H+\sigma^{\vee}.$
Since $H\subset\sigma^{\vee}$, $\lceil sp^{e}\rceil H+\sigma^{\vee}\subset sp^{e}H+\sigma^{\vee}$. Therefore, dividing through by $p^{e}$, we have that for $e\gg 0$,
$\frac{1}{p^{e}}\mathbf{c}+\mathbf{v}+\frac{\lceil tp^{e}+d\rceil}{p^{e}}H\subset\mathcal{B}+sH+\sigma^{\vee}.$
Now, $\mathcal{B}$ and $sH$ are each closed and bounded, hence compact, and therefore so is their sum $\mathcal{B}+sH$. Therefore, since $\sigma^{\vee}$ is closed, so is $\mathcal{B}+sH+\sigma^{\vee}$. Hence for any $\mathbf{h}\in H$,
$\mathbf{v}+t\mathbf{h}=\lim_{e\to\infty}\left(\frac{1}{p^{e}}\mathbf{c}+\mathbf{v}+\frac{\lceil tp^{e}+d\rceil}{p^{e}}\mathbf{h}\right)\in\mathcal{B}+sH+\sigma^{\vee}.$
@@
Given 6.1, suppose $k$ is of prime characteristic $p>0$, and let $t\in\mathbb{R}_{>0}$ and $x^{\mathbf{v}}\in R$ be a monomial.
@@
    If $\mathbf{v}+tH\subset\mathcal{B}+\sigma^{\vee}$, then $x^{\mathbf{v}}\in(I\overline{R})^{F\mathfrak{a}^{t}}$ and $x^{\mathbf{v}}\in(I\overline{R})^{*\mathfrak{a}^{t}}$.
@@
    If $\mathbf{v}+tH\subset\mathcal{B}+tH+\sigma^{\vee}$, then $x^{\mathbf{v}}\in{{}^{\mathfrak{a}^{t}}(I\overline{R})^{*}}$.
@@
    If $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ for some $s>t$, then $x^{\mathbf{v}}\in{}^{\mathfrak{a}^{t}}(I\overline{R})^{F}$.